OpenStudy (shaleiah):

What is the probability of rolling a sum of 5 on at least one of two rolls of a pair of number cubes?

OpenStudy (shaleiah):

a. 1/81 c. 2/9 b. 25/1296 d. 17/81

OpenStudy (shaleiah):

@563blackghost

OpenStudy (shaleiah):

@luckycoins888

OpenStudy (luckycoins888):

Wait up omg just wait xD

OpenStudy (luckycoins888):

my brain is still sleeping apparently i cant think for anything xD

OpenStudy (shaleiah):

Thank you so much!

OpenStudy (anonymous):

Made a mistake in the original posting. The odds of not rolling at least one 5, "no fives", in two rolls is: \[\left(\frac{32}{36}\right)^2 \]The odds of rolling at least one five is \[1-\left(\frac{32}{36}\right)^2=\frac{17}{81} \]